OpenStudy (anonymous):

5 square root of 6 times 2x square root 6

OpenStudy (anonymous):

\[5*\sqrt{6}*2x\sqrt{6}\] ?

OpenStudy (anonymous):

If so, when you have \[\sqrt{6}*\sqrt{6}\] They will become plain 6. And you'll get: 5*6*2x=?

OpenStudy (anonymous):

is that the answer?

Directrix (directrix):